OpenStudy (zuha):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y = –3x + 7 –2x + 6y = 3 Choose the correct answer: parallel perpendicular neither

OpenStudy (misty1212):

the slope of \(y=mx+b\) is \(m\) the slope of \(y=-3x+7\) is ?

OpenStudy (misty1212):

the slope of \(ax+by=c\) is \(-\frac{a}{b}\) the slope of \(-2x+6y=3\) is ?

OpenStudy (zuha):

niether

OpenStudy (zuha):

am I correct?

OpenStudy (anonymous):

You get the slopes 3 and -1/3. You see that they are not equal and therefore the lines aren't parallel, but what about perpendicular?

OpenStudy (anonymous):

niether

OpenStudy (zuha):

niether